If 70 mL of 0.01M HCL is added to 150 mL of 0.1M acetic acid, pH 5.0, what is the resulting pH? Identify the appropriate acid and conjugate base and determine their concentrations in the final solution. 


1
Expert's answer
2021-11-25T01:29:12-0500

Since we have initial pH of acetic acid before addition of HCl (pH=5.0)

We need to calculate the initial concentration of H3O+ ions


We know pH = 5.0 = -log [H3O+]

[H3O+] = 10-pH

= 10-5.0

= 0.00001 mol


What is the [H3O+] in the HCl added


Since HCl dissociates as;


HCl + H2O = H3O+ + Cl-


Moles of HCl = moles of H3O+


0.01mol were in 1000mL

How many moles will be in 70mL


"\dfrac{0.01mol x 70mL}{1000mL} = 0.0007mol"


Total [H3O+] = 0.00001mol + 0.0007mol = 0.00071mol


new pH = -Log(0.00071)

=3.15